OpenStudy (yacoub1993):

Solve the system by the method of your choice. Identify systems with no solution and systems with infinitely many solutions, using set notation to express their solution sets. x + y = -8 x - y = 14 A. {( 3, -11)} B. {( 3, 11)} C. {(x, y) | x + y = -8} D. ∅ <--------

OpenStudy (yacoub1993):

@Directrix

Directrix (directrix):

x + y = -8 x - y = 14 ----------- Add each term of the first equation to its like term in the second. Post what you get.

Directrix (directrix):

@Yacoub1993 Are you here with me?

Directrix (directrix):

Add x to x, then y to (-y) and then (-8) to 14

Directrix (directrix):

This which you marked above is not the correct answer. D. ∅ <-------- @Yacoub1993

OpenStudy (yacoub1993):

@Directrix 2x-y=6

OpenStudy (yacoub1993):

@Directrix i am having connection problem... sory

Directrix (directrix):

x + y = -8 x - y = 14 ----------- 2x + 0 = 6 2x = 6 x = ? @Yacoub1993

OpenStudy (yacoub1993):

x=3

Directrix (directrix):

And, to get y, 3 + y = -8 and y = -8 - 3 = ?

OpenStudy (yacoub1993):

x=3 and y = -11

Directrix (directrix):

Yes.

Directrix (directrix):

Option A, I think.

OpenStudy (yacoub1993):

thnxs

Directrix (directrix):

You are welcome.